A. Perform Horizontal SCI Analysis for 2020 uaing 2019 as base year.

1. What is the Peso Change and % Change if in 2020 the net sales is 1,687,500 while in 2019, the net sales is 1,350,000?

2. What is the Peso Change and % Change if in 2020 the Cost of Sales is 1,046,250 while in 2019, the Cost of Sales is 877,500?

3. What is the Peso Change and % Change if in 2020 the Selling Expenses is 270,000 while in 2019, the Selling Expenses is 162,000?

4. What is the Peso Change and % Change if in 2020 the Administrative Expenses is 135,000 while in 2019, the Administrative Expenses is 121,500?

5. What is the Peso Change and % Change if in 2020 the Interest Expense is 74,250 while in 2019, the Interest Expense is 54,000?

6. What is the Peso Change and % Change if in 2020 the Net Income is 162,000 while in 2019, the Net Income is 135,000?


1
Expert's answer
2021-10-08T17:11:40-0400

1. What is the Peso Change and % Change if in 2020 the net sales is 1,687,500 while in 2019, the net sales is 1,350,000?

Peso Change = 1,687,500 - 1,350,000 = +337,500

% Change = 1,687,500 / 1,350,000 * 100 - 100 = +25.00%

2. What is the Peso Change and % Change if in 2020 the Cost of Sales is 1,046,250 while in 2019, the Cost of Sales is 877,500?

Peso Change = 1,046,250 - 877,500 = +168,750

% Change = 1,046,250 / 877,500 * 100 - 100 = +19,23%

3. What is the Peso Change and % Change if in 2020 the Selling Expenses is 270,000 while in 2019, the Selling Expenses is 162,000?

Peso Change = 270,000 - 162,000 = +108,000

% Change = 270,000 / 162,000 * 100 - 100 = +66.67%

4. What is the Peso Change and % Change if in 2020 the Administrative Expenses is 135,000 while in 2019, the Administrative Expenses is 121,500?

Peso Change = 135,000 - 121,500 = +13,500

% Change = 135,000 / 121,500 * 100 - 100 = +11.11%

5. What is the Peso Change and % Change if in 2020 the Interest Expense is 74,250 while in 2019, the Interest Expense is 54,000?

Peso Change = 74,250 - 54,000 = +20,250

% Change = 74,250 / 54,000 * 100 - 100 = +37.50%

6. What is the Peso Change and % Change if in 2020 the Net Income is 162,000 while in 2019, the Net Income is 135,000?

Peso Change = 162,000 - 135,000 = +27,000

% Change = 162,000 - 135,000 * 100 - 100 = +20.00%